OpenStudy (anonymous):

A hamster ball has a radius of 21 centimeters. What is its volume written in terms of π? 12,348π cm3 6,946π cm3 588π cm3 9,261π cm3

OpenStudy (anonymous):

i believe it is \[V=\frac{4}{3}\pi r^3\]

OpenStudy (anonymous):

replace \(r\) by \(21\)

OpenStudy (anonymous):

Volume of sphere = 4/3 pi r^3 so , V = 4/3 pi 21^3